OpenStudy (anonymous):

A cylindrical water glass has a radius of 7 centimeters and a height of 14 centimeters. What is its volume rounded to the nearest hundredth?

OpenStudy (anonymous):

98

OpenStudy (anonymous):

\[V = \pi r^2 h\] \[= \pi*49*14\] \[Round that to the nearest hundreth\]

OpenStudy (anonymous):

@kumar2006 I dont think you're correct...

OpenStudy (anonymous):

2,155.13?

OpenStudy (anonymous):

Yes